# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v1

@@ Abstract
We consider the pricing of energy spread options for spot prices following an exponential Ornstein-Uhlenbeck process driven by a sum of independent multivariate variance gamma processes. Within this class of mean-reverting, infinite activity price processes, the Esscher transform is used to obtain an equivalent martingale measure. We focus on the weak variance alpha-gamma process and show that it is not closed under the Esscher transform. By deriving an analytic expression for the cumulant generating function of the innovation term, we then obtain a pricing formula for forwards and apply the FFT method of Hurd and Zhou to price spread options. Lastly, we demonstrate how the model should be both estimated on energy prices under the real world measure and calibrated on forward or call prices, and provide numerical results for the pricing of spread options.
@@ 1 Introduction
where $\boldsymbol{\Lambda}(t)$ is a deterministic seasonality function, and ${\bf X}$ is a Lévy-driven Ornstien-Uhlenbeck process (LDOUP) defined by
@@
where ${\bf Z}$ is the background driving Lévy process (BDLP), $\lambda>0$ is the speed of mean reversion, and the independent random vector ${\bf X}_{0}$ is the initial value. The Ornstien-Uhlenbeck part ensures mean reversion, while the Lévy part allows for jumps, and indeed the use of LDOUPs is widespread in energy price modeling. See for example BKMB ; benthbook ; BenthPir18 ; cf2005 ; sab2020 .
The BDLP ${\bf Z}$ we use herein is a sum of independent multivariate variance gamma (VG) processes with a focus on the weak variance alpha-gamma (WVAG) process by Buchmann, Lu and Madan BLM17a as a subclass. A multivariate VG process is a multivariate Brownian motion ${\bf B}$ subordinated, or time-changed, by a univariate gamma subordinator, and such processes originated with Madan and Seneta MaSe90 and have been studied in BKMS16 ; FS07 . While a sum of VG processes does not necessarily preserve such a time-change interpretation, the WVAG process being originally defined via weak subordination does. It turns out to be equivalent to a process ${\bf Z}$ characterized in the following way: consider a subordinator ${\bf T}$ which is a sum of $G_{k}$ representing the idiosyncratic time-change for the $k$th price, and $\boldsymbol{\alpha}G_{0}$ representing a common time change for all prices, where $G_{0},\dots,G_{n}$ are independent gamma subordinators and $\boldsymbol{\alpha}\in[0,\infty)^{n}$, then when ${\bf T}(t)$ jumps by $\Delta{\bf T}(t)$, ${\bf Z}(t)$ jumps by ${\bf B}(\Delta{\bf T}(t))$ in distribution, rather than pathwisely for traditional subordination. Consequently, the WVAG process, using weak subordination, preserves the essence of traditional subordination in a distributional or weak sense.
@@
For the pricing of energy derivatives, we use the Esscher transform gs94 to obtain an equivalent martingale measure (EMM) or Esscher measure $\mathbb{Q}_{{\bf h}}$ indexed by the market price of risk ${\bf h}$. This is a common approach for Lévy-based models, in which the market is incomplete due to the presence of jumps. Energy prices such as electricity are nontraded assets in the sense that term is used in standard arbitrage theory since electricity bought or sold on the market is instantly consumed and cannot be held to construct a replicating portfolio for the valuation of derivatives. Consequently, it is not assumed that the discounted energy price process is a $\mathbb{Q}_{{\bf h}}$-martingale, a condition that would ordinarily determine a unique ${\bf h}$ and hence an EMM $\mathbb{Q}_{{\bf h}}$ in an equity market. Instead, in an energy market, ${\bf h}$ must be obtained by calibration, such as to observed forward and call prices.
@@
For the WVAG BDLP, we provide numerical results under both the true model and the fitted model. Under the true model, the spread option price is computed via the analytical formula for the cgf of the log return with the FFT method, and is compared to the Monte Carlo method, showing close agreement. We also provide an end-to-end implementation of the spread option pricing in a simulation study by fitting the model. Our model allows the seasonality and LDOUP parameters to be estimated under the real world measure $\mathbb{P}$, specifically using the MLE method by Lu Lu21 combined with the analytical cgf of the log return, while the market price of risk ${\bf h}$ is estimated under the Esscher measure $\mathbb{Q}_{{\bf h}}$, in the sense of calibrating the model to observed forward and call prices. Thus, we can coherently leverage both data on the $\mathbb{P}$-dynamics of the underlying and its derivative prices for model fitting, in contrast to the practice of fully calibrating the model under the risk-neutral measure for pricing energy derivatives gs22 ; sab2020 , and more commonly equity derivatives MCC98 ; sco . Finally, the fitted model can be used to price spread options over a panel of strikes. We find in our example that calibrating to calls is better than calibrating to forwards, and demonstrate the extent to which errors in estimating the model parameters propagate to pricing errors.
@@ 2 Preliminaries
We use the convention that ${\bf x}=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ represents a column vector. Let ${\boldsymbol{\Sigma}}=(\Sigma_{ij})_{ij=1}^{n}\in\mathbb{R}^{n\times n}$ denote an $n\times n$ real matrix. For a multi-index $\alpha$, define $[\cdot]_{\alpha}$ as extracting the $\alpha$th element. For $A\subseteq\mathbb{R}^{n}$, let ${\bf 1}_{A}$ be its indicator function and $A_{*}:=A\backslash\{{\bf 0}\}$. Let $\mathbb{D}:=\{{\bf x}\in\mathbb{R}^{n}:\|{\bf x}\|\leq 1\}$. For ${\bf x},{\bf y}\in\mathbb{R}^{n}$, define the Euclidean inner product $\langle{\bf x},{\bf y}\rangle:={\bf x}^{\prime}{\bf y}$, and $\|{\bf x}\|_{{\boldsymbol{\Sigma}}}:=\sqrt{\langle{\bf x},{\boldsymbol{\Sigma}}{\bf x}\rangle}$, which is a norm when ${\boldsymbol{\Sigma}}$ is positive definite. Let ${\bf e}_{k}\in\mathbb{R}^{n}$ be the $k$th standard basis vector. Let $I:[0,\infty)\to[0,\infty)$ be the identity function. The functions $\exp$ and $\log$ are understood to act componentwise on vectors so, for example, $e^{\bf x}=(e^{x_{1}},\dots,e^{x_{n}})$.
@@
We write a process ${\bf X}$ in terms of its marginal components and time marginals by ${\bf X}=(X_{1},\dots,X_{n})=({\bf X}(t))_{t\geq 0}$. Let =$\stackrel{{\scriptstyle D}}{{=}}$ denote equality in distribution or law. Let $\kappa_{{\bf X}}$, $\Psi_{{\bf X}}$ and $\Phi_{{\bf X}}$ denote respectively the cumulant generating function (cgf), characteristic exponent and characteristic function of the random vector ${\bf X}$, or of ${\bf X}(1)$ if ${\bf X}$ is a process, under $\mathbb{P}$. Specifically, for a random vector ${\bf X}$, these are defined by
@@
Let $\kappa_{{\bf X}}^{\mathbb{Q}}$, $\Phi_{{\bf X}}^{\mathbb{Q}}$ and $\Psi_{{\bf X}}^{\mathbb{Q}}$ denote the respective functions under $\mathbb{Q}$. Let ${\bf B}\sim BM^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ be an $n$-dimensional Brownian motion with drift $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and covariance ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$. Let $\circ$ denote the componentwise composition. For ${\bf t}=(t_{1},\dots,t_{n})\in[0,\infty)^{n}$, define
@@
which are the mean vector and covariance matrix of ${\bf B}\circ{\bf t}=(B_{1}(t_{1}),\dots,B_{n}(t_{n}))$.
@@ 2.1 Variance gamma and related process
We begin with a summary of relevant definitions and results from the studies BKMS16 ; BLM17a . Let $\Gamma_{S}(a,b)$ denote a univariate gamma subordinator with shape $a>0$ and rate $b>0$. An $n$-dimensional process ${\bf V}\sim VG^{n}(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is a variance gamma (VG) process if
${\bf V}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf B}\circ(G,\dots,G),$
@@
${\cal D}_{{\bf V}}$ :=$:=\{\boldsymbol{\theta}\in\mathbb{R}^{n}\,:\,K_{\boldsymbol{\theta}}>0\}.$ (2.3)
@@
We define ${\bf Z}\sim{VGC}^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a VG convolution (VGC) or sum of VG processes if
@@
Let $n\geq 2$, $a>0$, $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{n})\in(0,1/a)^{n}$, $\beta_{k}:=(1-a\alpha_{k})/{\alpha_{k}}$, $k=1,\dots,n$, and $\boldsymbol{\eta}\in\mathbb{R}^{n}$. We define ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ as a weak variance alpha-gamma (WVAG) process if
@@
where ${\bf V}_{0}\sim VG^{n}(a,a\boldsymbol{\mu}\diamond\boldsymbol{\alpha},a{\boldsymbol{\Sigma}}\diamond\boldsymbol{\alpha})$ and $V_{k}\sim VG^{1}(\beta_{k},\allowbreak\alpha_{k}\beta_{k}\mu_{k},\alpha_{k}\beta_{k}\Sigma_{kk})$, $k=1,\dots,n$, are independent. This representation allows us to see that the WVAG process is a sum of VG processes with ${\bf Z}\sim VGC^{n}_{n+1}({\bf b}_{S},\boldsymbol{\mu}_{S},{\boldsymbol{\Sigma}}_{S},\boldsymbol{\eta})$, where
${\bf b}_{S}$ $=(a,\beta_{1},\dots,\beta_{n}),$ (2.6)
$\boldsymbol{\mu}_{S}$ $=(a\boldsymbol{\mu}\diamond\boldsymbol{\alpha},\alpha_{1}\beta_{1}\mu_{1}{\bf e}_{1},\dots,\alpha_{n}\beta_{n}\mu_{n}{\bf e}_{n}),$ (2.7)
${\boldsymbol{\Sigma}}_{S}$ $=(a{\boldsymbol{\Sigma}}\diamond\boldsymbol{\alpha},\alpha_{1}\beta_{1}\Sigma_{11}{\bf e}_{1}{\bf e}_{1}^{\prime},\dots,\alpha_{n}\beta_{n}\Sigma_{nn}{\bf e}_{n}{\bf e}_{n}^{\prime}),$ (2.8)
@@ 2.2 Lévy-Driven Ornstein-Uhlenbeck Processes
Let ${\bf Z}\sim L^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}},{\cal Z})$ denote an $n$-dimensional Lévy process with characteristic triplet $(\boldsymbol{\mu},{\boldsymbol{\Sigma}},{\cal Z})$, where $\boldsymbol{\mu}\in\mathbb{R}^{n}$, ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$ is a covariance matrix, and ${\cal Z}$ is a Lévy measure. With reference to Lu21 ; mas04 ; sato99 ; saya , we recall that a Lévy-driven Ornstein-Uhlenbeck ${\bf X}\sim{OU}\text{-}{\bf Z}(\lambda)$ is defined by the SDE (1.2) for $t\geq 0$, where ${\bf Z}\sim L^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}},{\cal Z})$ is the BDLP, $\lambda>0$, and ${\bf X}_{0}$ is independent of ${\bf Z}$. Note that there is no loss in generality in having the drift term $-\lambda{\bf X}(t){\rm d}t$ instead of $\lambda(\overline{\boldsymbol{\mu}}-{\bf X}(t)){\rm d}t$ for some $\overline{\boldsymbol{\mu}}\in\mathbb{R}^{n}$ since the stationary mean of ${\bf X}$ is controlled by $\mathbb{E}[{\bf Z}(1)]$. There is also no loss in generality in having the Lévy term ${\rm d}{\bf Z}(\lambda t)$ instead of ${\rm d}\widetilde{\bf Z}(t)$ for some general Lévy process $\widetilde{\bf Z}\sim L^{n}(\widetilde{\boldsymbol{\mu}},\widetilde{\boldsymbol{\Sigma}},\widetilde{\cal Z})$ since we can write $\widetilde{\bf Z}\stackrel{{\scriptstyle D}}{{=}}{\bf Z}\circ(\lambda I)$, where ${\bf Z}\sim L^{n}(\widetilde{\boldsymbol{\mu}}/\lambda,\widetilde{\boldsymbol{\Sigma}}/\lambda,\widetilde{\cal Z}/\lambda)$.
@@ 3 Analytical formula for cgf of the innovation term
This formula, which will later be applied under both $\mathbb{P}$ and $\mathbb{Q}_{{\bf h}}$, has 3 applications in the pricing of energy derivatives. First, in the univariate and bivariate setting under $\mathbb{P}$, it is used in the likelihood function to estimate the LDOUP parameters (see Section 5.3). Second, in the univariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price forwards (see Section 4.4.1), calls (see Section 4.4.2), and calibrate the model (see Section 5.4). Third, in the bivariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price spread options (see Sections 4.4.3 and 5.5).
To begin, for a general ${\bf X}\sim OU\text{-}{\bf Z}(\lambda)$, $\lambda>0$, we give a lemma to deal with finiteness of the cgf, which is a version of (bsb04, , Lemma 4.1) adapted to our setting with a slightly different proof.
@@
${\cal Z}^{*}(B)=\int_{\mathbb{R}^{n}_{*}}\int_{0}^{t}\lambda{\bf 1}_{e^{\lambda t}B}(e^{-\lambda s}{\bf z})\,{\rm d}s{\cal Z}({\rm d}{\bf z}),\quad B\in{\cal B}(\mathbb{R}^{n}_{*}),$
@@
for all $u\in[0,\lambda t]$. Next, by a version of the dominated convergence theorem (see (Ba01, , Lemma 16.1)) using ${\bf z}\mapsto 1+e^{\langle e^{\lambda t}\boldsymbol{\theta},{\bf z}\rangle}$ as the dominating function and the definition of the Lévy measure for ${\cal Z}$, it follows that $f$ is continuous on $[0,\lambda T]$. Thus, $I<\infty$, so that $\kappa_{{\bf Z}^{*}(t)}(\boldsymbol{\theta})$ is finite for all $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(t)}$ by the exponential moment theorem.
@@
where $\1z\in\mathbb{C}\backslash[1,\infty)$. Provided that $z\leq 1$, $\operatorname{Li}_{2}(z)\in\mathbb{R}$ with $\operatorname{Li}_{2}(1)=\pi^{2}/6$. Now we give the analytical formula for the cgf of ${\bf V}^{*}(t)$ using a method inspired by sab2020 .
@@
We now check that the above expressions are defined. Firstly, according to Lemma 3.1, $M(e^{s})>0$ for $s\in[0,\lambda t]$, or equivalently $M(u)>0$ for $u\in[1,c]$, so (3.4) is defined. The coefficient of $u^{2}$ in $M(u)$ satisfies $AB=\|\boldsymbol{\theta}\|_{{\boldsymbol{\Sigma}}}^{2}/(2b)>0$. Clearly $A>0$, so we must have $B>0$. Now either both $1-Au$ and $1+Bu$ are strictly positive for all $u\in[1,c]$, or strictly negative. However, $01+Bu<$ 0 is impossible since $u<-1/B<0$ contradicts $u\geq 1$. Thus, we must have $01-Au>$ 0 and $01+Bu>$ 0 for all $u\in[1,c]$, so (3.5) is defined. Observe that $-cB<B\leq A<cA<1$, where the last inequality is implied by $01-Au>$ 0 on $u\in[1,c]$. Thus, we have ensured (3.6) can be evaluated using $\operatorname{Li}_{2}$ on the subset of the domain where the function is real.
@@
This is why Sabino’s formula has $|\theta|$ in (3.8) replaced with5 $\theta$ but is otherwise the same. Numerically, this causes the arguments of $\operatorname{Li}_{2}$ to be different despite the value of $\kappa_{V^{*}(t)}(\theta)$ being the same.
@@
which raises the question of whether ${\bf V}$ can be written as the difference of two other Lévy processes ${\bf G}^{\pm}$ with characteristic exponents $\Psi_{{\bf G}^{\pm}}(\boldsymbol{\theta})$. Unfortunately, the answer is no as $\Psi_{{\bf G}^{\pm}}(\boldsymbol{\theta})$ is not a valid characteristic exponent. If it were, then its first marginal component $G_{1}^{\pm}$ would have characteristic exponent
$\Psi_{G_{1}^{\pm}}(\theta_{1})=-b\log\mathopen{}\left(1-{\rm i}\frac{\sqrt{\mu_{1}^{2}+2b\Sigma_{11}}|\theta_{1}|\pm\mu_{1}\theta_{1}}{2b}\right)\mathclose{},\quad\theta_{1}\in\mathbb{R},$

### v2

@@ Abstract
We consider the pricing of energy spread options for spot prices following an exponential Ornstein-Uhlenbeck process driven by a sum of independent multivariate variance gamma processes, which gives rise to mean-reverting, infinite activity price dynamics. Within this class of driving processes, the Esscher transform is used to obtain an equivalent martingale measure with a focus on the weak variance alpha-gamma process. By deriving an analytic expression for the cumulant generating function of the innovation term, we obtain a pricing formula for forwards and apply the FFT method of Hurd and Zhou to price spread options. Lastly, we demonstrate how the model should be both estimated on energy prices under the real world measure and calibrated on forward or call prices, and provide numerical results for the pricing of spread options.
@@ 1 Introduction
where $\boldsymbol{\Lambda}(t)$ is a deterministic seasonality function, and $\bf X$} is a Lévy-driven Ornstein-Uhlenbeck process (LDOUP) defined by
@@
where $\bf Z$} is the background driving Lévy process (BDLP), $\lambda>0$ is the speed of mean reversion, and the independent random vector ${\bf X}_{0}$ is the initial value. The Ornstein-Uhlenbeck part ensures mean reversion, while the Lévy part allows for jumps, and indeed the use of LDOUPs is widespread in energy price modeling. See for example BKMB ; benthbook ; BenthPir18 ; cf2005 ; sab2020 ; sab2023 .
The BDLP $\bf Z$} we use herein is a sum of independent multivariate variance gamma (VG) processes with a focus on the weak variance alpha-gamma (WVAG) process by Buchmann, Lu and Madan BLM17a as a subclass. A multivariate VG process is a multivariate Brownian motion $\bf B$} subordinated, or time-changed, by a univariate gamma subordinator, and such processes originated with Madan and Seneta MaSe90 and have been studied in BKMS16 ; FS07 . While a sum of VG processes does not necessarily preserve such a time-change interpretation, the WVAG process being originally defined via weak subordination does. It turns out to be equivalent to a process $\bf Z$} characterized in the following way: consider a subordinator $\bf T$} which is a sum of $G_{k}$ representing the idiosyncratic time-change for the $k$th price, and $\boldsymbol{\alpha}G_{0}$ representing a common time change for all prices, where $G_{0},\dots,G_{n}$ are independent gamma subordinators and $\boldsymbol{\alpha}\in[0,\infty)^{n}$, then when ${\bf T}(t)$ jumps by $\Delta{\bf T}(t)$, ${\bf Z}(t)$ jumps by the componentwise composition ${\bf B}(\Delta{\bf T}(t))$ in distribution, rather than pathwisely for traditional subordination. Consequently, the WVAG process, using weak subordination, preserves the essence of traditional subordination in a distributional or weak sense.
@@
For the pricing of energy derivatives, we use the Esscher transform gs94 to obtain an equivalent martingale measure (EMM) or Esscher measure $\mathbb{Q}_{{\bf h}}$ indexed by the market price of risk (MPR) $\bf h$}. This is a common approach for Lévy-based models, in which the market is incomplete due to the presence of jumps. Energy prices such as electricity are nontraded assets in the sense that term is used in standard arbitrage theory since electricity bought or sold on the market is instantly consumed and cannot be held to construct a replicating portfolio for the valuation of derivatives. Consequently, it is not assumed that the discounted energy price process is a $\mathbb{Q}_{{\bf h}}$-martingale, a condition in an equity market that would ordinarily determine a unique $\bf h$}, and hence an EMM $\mathbb{Q}_{{\bf h}}$. Instead, in an energy market, $\bf h$} must be obtained by calibration, such as to observed forward or call prices.
@@
For the WVAG BDLP, we provide numerical results under both the true model and the fitted model. Under the true model, the spread option price is computed via the analytical formula for the cgf of the log return with the FFT method, and is compared to the Monte Carlo method, showing close agreement. We also provide an end-to-end implementation of the spread option pricing in a simulation study by fitting the model. Our model allows the seasonality and LDOUP parameters to be estimated under the real world measure $\mathbb{P}$, specifically using the MLE method by Lu Lu21 combined with the analytical formula for the cgf of the innovation term, while the MPR $\bf h$} is estimated under the Esscher measure $\mathbb{Q}_{{\bf h}}$, in the sense of calibrating the model to observed forward and call prices. Thus, we can coherently leverage both data on the $\mathbb{P}$-dynamics of the underlying and its derivative prices for model fitting, in contrast to the practice of fully calibrating the model under the risk-neutral measure for pricing energy derivatives gs22 ; sab2020 , and more commonly equity derivatives MCC98 ; sco . Finally, the fitted model can be used to price spread options over a panel of strikes. We find in our example that calibrating to calls is better than calibrating to forwards, and demonstrate the extent to which errors in estimating the model parameters propagate to pricing errors.
@@ 2 Preliminaries
[…]
@@
We write a process $\bf X$} in terms of its marginal components and time marginals by ${\bf X}=(X_{1},\dots,X_{n})=({\bf X}(t))_{t\geq 0}$. Let $\stackrel{{\scriptstyle D}}{{=}}$ denote equality in distribution or law. Let $\kappa_{{\bf X}}$, $\Psi_{{\bf X}}$ and $\Phi_{{\bf X}}$ denote respectively the cumulant generating function (cgf), characteristic exponent and characteristic function of the random vector $\bf X$}, or of ${\bf X}(1)$ if $\bf X$} is a process, under $\mathbb{P}$. Specifically, for a random vector $\bf X$}, these are defined by
@@
Let $\kappa_{{\bf X}}^{\mathbb{Q}}$, $\Phi_{{\bf X}}^{\mathbb{Q}}$ and $\Psi_{{\bf X}}^{\mathbb{Q}}$ denote the respective functions under $\mathbb{Q}$. Let ${\bf B}\sim BM^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ be an $n$-dimensional Brownian motion with drift $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and covariance ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$. For ${\bf t}=(t_{1},\dots,t_{n})\in[0,\infty)^{n}$, define
@@
which are the mean vector and covariance matrix of the componentwise composition ${\bf B}({\bf t})$.
@@ 2.1 Variance gamma and related process
We begin with a summary of relevant definitions and results from the studies BKMS16 ; BLM17a . Let $\Gamma_{S}(a,b)$ denote a univariate gamma subordinator with shape $a>0$ and rate $b>0$. An $n$-dimensional ${\bf V}\sim VG^{n}(b,\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is a variance gamma (VG) process is defined by
${\bf V}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf B}(G),$
@@
${\cal D}_{{\bf V}}$ $:=\{\boldsymbol{\theta}\in\mathbb{R}^{n}\,:\,K_{\boldsymbol{\theta}}>0\}.$ (2.3)
@@
A VG convolution (VGC) or sum of VG processes ${\bf Z}\sim{VGC}^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is defined by
@@
Let $n\geq 2$, $a>0$, $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{n})\in(0,1/a)^{n}$, $\beta_{k}:=(1-a\alpha_{k})/{\alpha_{k}}$, $k=1,\dots,n$, and $\boldsymbol{\eta}\in\mathbb{R}^{n}$. A weak variance alpha-gamma (WVAG) process ${\bf Z}\sim WVAG^{n}(a,\boldsymbol{\alpha},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is defined by
@@
where ${\bf V}_{0}\sim VG^{n}(a,a\boldsymbol{\mu}\diamond\boldsymbol{\alpha},a{\boldsymbol{\Sigma}}\diamond\boldsymbol{\alpha})$ and $V_{k}\sim VG^{1}(\beta_{k},\allowbreak\alpha_{k}\beta_{k}\mu_{k},\alpha_{k}\beta_{k}\Sigma_{kk})$, $k=1,\dots,n$, are independent. This representation allows us to see that the WVAG process is a sum of VG processes with ${\bf Z}\sim VGC^{n}_{n+1}({\bf b}_{C},\boldsymbol{\mu}_{C},{\boldsymbol{\Sigma}}_{C},\boldsymbol{\eta})$, where
${\bf b}_{C}$ $=(a,\beta_{1},\dots,\beta_{n}),$ (2.6)
$\boldsymbol{\mu}_{C}$ $=(a\boldsymbol{\mu}\diamond\boldsymbol{\alpha},\alpha_{1}\beta_{1}\mu_{1}{\bf e}_{1},\dots,\alpha_{n}\beta_{n}\mu_{n}{\bf e}_{n}),$ (2.7)
${\boldsymbol{\Sigma}}_{C}$ $=(a{\boldsymbol{\Sigma}}\diamond\boldsymbol{\alpha},\alpha_{1}\beta_{1}\Sigma_{11}{\bf e}_{1}{\bf e}_{1}^{\prime},\dots,\alpha_{n}\beta_{n}\Sigma_{nn}{\bf e}_{n}{\bf e}_{n}^{\prime}),$ (2.8)
@@ 2.2 Lévy-Driven Ornstein-Uhlenbeck Processes
[…]
@@ 3 Analytical formula for cgf of the innovation term
This formula, which is stated in Theorem 3.3 and will later be applied under both $\mathbb{P}$ and $\mathbb{Q}_{{\bf h}}$, has 3 applications in the pricing of energy derivatives. Firstly, in the univariate and bivariate setting under $\mathbb{P}$, it is used in the likelihood function to estimate the LDOUP parameters (see Section 5.3). Secondly, in the univariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price forwards (see Section 4.4.1), calls (see Section 4.4.2), and calibrate the model (see Section 5.4). Thirdly, in the bivariate setting under $\mathbb{Q}_{{\bf h}}$, it is used to price spread options (see Sections 4.4.3 and 5.5).
To begin, for a general ${\bf X}\sim OU\text{-}{\bf Z}(\lambda)$, $\lambda>0$, we give a lemma to deal with the finiteness of the cgf, which is a version of (bsb04, , Lemma 4.1) adapted to our setting with a slightly different proof.
@@
${\cal Z}^{*}(B)=\int_{\mathbb{R}^{n}\backslash\{{\bf 0}\}}\int_{0}^{t}\lambda{\bf 1}_{e^{\lambda t}B}(e^{-\lambda s}{\bf z})\,{\rm d}s{\cal Z}({\rm d}{\bf z}),\quad B\in{\cal B}(\mathbb{R}^{n}_{*}),$
@@
for all $u\in[0,\lambda t]$. Next, by a version of the dominated convergence theorem (see (Ba01, , Lemma 16.1)) using ${\bf z}\mapsto 1+e^{\langle e^{\lambda t}\boldsymbol{\theta},{\bf z}\rangle}$ as the dominating function and the definition of the Lévy measure for $\cal Z$}, it follows that $f$ is continuous on $0,\lambda T]$. Thus, $I<\infty$, so that $\kappa_{{\bf Z}^{*}(t)}(\boldsymbol{\theta})$ is finite for all $\boldsymbol{\theta}\in{\cal D}_{{\bf Z}^{*}(t)}$ by the exponential moment theorem.
@@
where $z\in\mathbb{C}\backslash[1,\infty)$. Provided that $z\leq 1$, $\operatorname{Li}_{2}(z)\in\mathbb{R}$ with $\operatorname{Li}_{2}(1)=\pi^{2}/6$. Now we give the analytical formula for the cgf of ${\bf V}^{*}(t)$ using a method inspired by sab2020 .
@@
We now check that the above expressions are defined. Firstly, according to Lemma 3.1, $M(e^{s})>0$ for $s\in[0,\lambda t]$, or equivalently $M(u)>0$ for $u\in[1,c]$, so (3.4) is defined. The coefficient of $u^{2}$ in $M(u)$ satisfies $AB=\|\boldsymbol{\theta}\|_{{\boldsymbol{\Sigma}}}^{2}/(2b)>0$. Clearly $A>0$, so we must have $B>0$. Now either both 1Au1-Au and $1+Bu$ are strictly positive for all $u\in[1,c]$, or strictly negative. However, $1+Bu<0$ is impossible since u<1/B<0u<-1/B<0 contradicts $u\geq 1$. Thus, we must have 1Au>01-Au>0 and $1+Bu>0$ for all $u\in[1,c]$, so (3.5) is defined. Observe that $cB<B\leq A<cA<1$, where the last inequality is implied by 1Au>01-Au>0 on $u\in[1,c]$. Thus, we have ensured (3.6) can be evaluated using $\operatorname{Li}_{2}$ on the subset of the domain where the function is real.
@@
This is why Sabino’s formula has $|\theta|$ in (3.8) replaced with $\theta$ but is otherwise the same. Numerically, this causes the arguments of $\operatorname{Li}_{2}$ to be different despite the value of $\kappa_{V^{*}(t)}(\theta)$ being the same.
@@
which raises the question of whether $\bf V$} can be written as the difference of two other Lévy processes $\bf G$$^{\pm}$ with characteristic exponents $\Psi_{{\bf G}^{\pm}}(\boldsymbol{\theta})$. Unfortunately, the answer is no as $\Psi_{{\bf G}^{\pm}}(\boldsymbol{\theta})$ is not a valid characteristic exponent. If it were, then the first marginal component $G_{1}^{+}$ would have characteristic exponent
$\Psi_{G_{1}^{+}}(\theta_{1})=-b\log\mathopen{}\left(1-{\rm i}\frac{\sqrt{\mu_{1}^{2}+2b\Sigma_{11}}|\theta_{1}|+\mu_{1}\theta_{1}}{2b}\right)\mathclose{},\quad\theta_{1}\in\mathbb{R},$